OpenStudy (anonymous):

A particle moves along the x-axis so that at time t its position is given by s(t) = (t + 3)(t −1)3, t > 0. For what values of t is the velocity of the particle increasing?

OpenStudy (danjs):

the change in position function s(t) w.r.t. time is the Velocity function , V(t) V(t) = s ' (t)

OpenStudy (anonymous):

Huh?

OpenStudy (danjs):

have you done derivatives,

OpenStudy (anonymous):

yep

OpenStudy (danjs):

take the derivative of the position function, that will be the velocity function... if you take a second derivative, that is then the acceleration function

OpenStudy (danjs):

Velocity will be increasing when you have a net acceleration in the same direction you moving

OpenStudy (anonymous):

So when t>1?

OpenStudy (danjs):

it is like the increasing /decreasing first derivative test from calc 1,

OpenStudy (danjs):

is that a cube on the (t - 1)^3 ?

OpenStudy (anonymous):

Yes.

OpenStudy (danjs):

s(t) = (t + 3)(t −1)^3 v(t) = 4*(t - 1)^2 * (t+2) a(t) = 12*(t - 1)*(t + 1) position, velocity, acceleration functions

OpenStudy (danjs):

They ask when v(t) is increasing in value. you can look at the acceleration function a(t) to decide this, since accel is the rate of change of velocity.

OpenStudy (anonymous):

So the answer isn't t>1?

OpenStudy (anonymous):

My choices are 0 < t < 1 t > 1 t > 0 The velocity is never increasing

OpenStudy (danjs):

oh i dont know yet

OpenStudy (danjs):

yeah looks like + acceleration and increased velocity when t is larger than 1

OpenStudy (danjs):

0 to 1, is negative values

OpenStudy (anonymous):

Thank you. Can you also help me with another question?

OpenStudy (danjs):

SURE BRB

OpenStudy (anonymous):

Okie. A balloon rises at the rate of 8 feet per second from a point on the ground 12 feet from an observer. To 2 decimal places in radians per second, find the rate of change of the angle of elevation when the balloon is 9 feet above the ground. Type your answer in the space below. If your answer is a number less than 1, place a leading "0" before the decimal point (ex: 0.35). Whenever you are ready. cx

OpenStudy (empty):

What"s your attempt?

OpenStudy (anonymous):

I have non. In all honesty, I'm just trying to finish a pretest as fast as I can with a not so shabby grade, so I can get back to my ACT prep.

OpenStudy (anonymous):

none

OpenStudy (danjs):

yeah this related rate prob..

OpenStudy (empty):

so youre just getting dan to answer your question for you? smh

OpenStudy (empty):

It just looks pretty disrespectful to string him along like this, pretending to be following along while he gives you the answer when he's doing this for free out of the goodness of his own heart and all.

OpenStudy (danjs):



OpenStudy (anonymous):

I'm also listening and taking into account what he has to say. You make it sound like I;m just about the answer, I do want to see how it is done because I need it, but at the moment I prefer a fast explained and concise answer for future reference. I apologize that my ACT is a bit more important to me. @Empty

OpenStudy (anonymous):

Okay, why is 8 dh/dt?

OpenStudy (danjs):

rises at 8ft per second, that is the rate of change of height h, dh/dt

OpenStudy (anonymous):

Ohhh, so what would be the next step? The picture helps a lot by the way, cx

OpenStudy (danjs):

you need something to relate the angle and the height, it is a right triangle

OpenStudy (anonymous):

Do I just solve or a when h is 9?

OpenStudy (danjs):

THis is dynamic, the height is changing and the angle is changing at the same time always, they ask for how fast the angle is changing at that instant in time, when h=9...

OpenStudy (danjs):

you need some relationship between the angle and the height then take the derivative with respect to time of everything, so you then have a relationship between the rates of change of height and angle...

OpenStudy (anonymous):

I'm confused.

OpenStudy (danjs):

maybe the tangent will work... tan(a) = h / 12

OpenStudy (danjs):

relationship between the angle and the height

OpenStudy (anonymous):

DO I use 9 for h?

OpenStudy (danjs):

, that will tell you the angle 'a' at that instant... we need to get the rate the angle is changing

OpenStudy (danjs):

take the derivative of both sides of the function with respect to time

OpenStudy (anonymous):

Of the tangent equation?

OpenStudy (danjs):

yes, need to use the chain rule or implicit differentiating derivative of tan(a) will be first derivative with respect to angle a, then da/dt to give you d/dt of the function

OpenStudy (danjs):

tan(a) = h/12 \[\sec^2(a) * \frac{ da }{ dt } = \frac{ 1 }{ 12 }*\frac{ dh }{ dt }\]

OpenStudy (anonymous):

I did it through and got 0.425. Does that sound correct>

OpenStudy (danjs):

i dont know

OpenStudy (danjs):

\[\frac{ da }{ dt }=\frac{ 1 }{ 12 }*\cos^2(a)*\frac{ dh }{ dt }\] so tha tis the rate the angle a changes in terms of the rate h is changing

OpenStudy (danjs):

dh/dt is given at 8 ft/s the angle 'a' needed is found from the instant when h=9 from the tangent tan(a)=9/12 a = 0.6435 that is it, calculate the da/dt

OpenStudy (danjs):

yeah da/dt = 0.43 rad / s when the baloon is at h=9 ft

OpenStudy (danjs):

unless i f'ed up

OpenStudy (anonymous):

YAY!!!

OpenStudy (anonymous):

Do you know anything about Rolles Theorem?

OpenStudy (danjs):

all those probs are similar, need a relationship between variables, and the problem usually has it for you... derivative of both sides, and maybe some substitutions

OpenStudy (danjs):

maybe, is that one of those existence theorems... like intermediate value, mean value

OpenStudy (anonymous):

Yep.

OpenStudy (danjs):

i remember the name, i forget what it says

OpenStudy (anonymous):

Thats okay. Thank you for all the help. Have a great day and life. cx

OpenStudy (danjs):

k goodluck